OpenStudy (anonymous):

Point G is located at (3, -1) and point H is located at (-2, 3). Find the point that is 2/3 the distance from point G to point H. (5 points) Select one: a. (0.33, -1.67) b. (-0.33, 1.67) c. (-0.5, -1) d. (6.33, -2.67)

OpenStudy (anonymous):

Go 2/3 the distance in x from G to H and 2/3 the distance in y from G to H. These are the new coordinates.

OpenStudy (anonymous):

Parametrize the line from G to H as f(t) =G + t (H - G)={3 - 5 t, -1 + 4 t} t between 0 and 1 Notice the f[0)=G f[1]= H f(1/2)= Midpoint from G to H You need \[ f(2/3)=\left(-\frac{1}{3},\frac{5}{3}\right)=(-0.333333, 1.66667) \]